# HippoRAG: Neurobiologically Inspired Long-Term Memory for Large Language Models

### v1

@@ Appendix G Cost and Efficiency Comparison
Although offline indexing time and costs are higher for HippoRAG than IRCoT—around $10$ times slower and \$$15$ more expensive for every $10,000$ passages  [^7] To speed up indexing, we use $10$ threads querying gpt-3.5-turbo-1106 through the OpenAI API in parallel. At the time of writing, the cost of the API is \$$1$ for a million input tokens and \$$2$ for a million output tokens., these costs can be dramatically reduced by leveraging open source LLMs. As shown in our ablation study in Table 5 Llama-3-8B-Instruct [1] performs similarly to GPT-3.5 Turbo without any specific prompt tuning. As seen in Table 16, this 8B Llama-3 model is about half the cost as GPT-3.5 Turbo when using an LLM deploying service like Together AI  [^8] https://www.together.ai/products#inference. Additionally, since these costs could be even further reduced by locally deploying this model, the barriers for using HippoRAG at scale could be well within the computational budget of many organizations. Finally, we note that even if LLM generation cost drops significantly, the online retrieval efficiency gains discussed above remain intact given that the number of tokens required for IRCoT vs. HippoRAG stay constant and LLM use is likely to also remain the system’s main computational bottleneck.
Table 16: Average cost and latency measurements for offline indexing using GPT-3.5 Turbo and Llama-3 (8B and 70B) through the TogetherAI API on $10,000$ passages.
@@
Llama-3-8B-Instruct API Cost (\$) 00 00 $3$
Time (minutes) $7$ $7$ $45$
Llama-3-70B-Instruct API Cost (\$) 00 00 $7$
Time (minutes) $7$ $7$ $100$
@@ Appendix H Implementation Details & Compute Requirements
In terms of compute requirements, most of our compute requirements are unfortunately not disclosed by the OpenAI and TogetherAI APIs. We run ColBERTv2 and Contriever for indexing and retrieval we use 4 NVIDIA RTX A6000 GPUs with 48GB of memory. Finally, we used 2 AMD EPYC 7513 32-Core Processors to run the Personalized PageRank algorithm.

### v3

@@ F.3 OpenIE Limitations
F.4 OpenIE Document Length Analysis
Finally, we present a small-scale intrinsic experiment to help us understand the robustness of our OpenIE methods to increasing passage length. The length-dependent evaluation results in Table 16, show that GPT-3.5-Turbo OpenIE results deteriorate substantially when extracting from longer instead of shorter passages. This is likely due to a higher sentence and paragraph complexity for longer passages which leads to lower quality extraction. More work is needed to address this limitation since further chunking would only create other issues due to sentence interdependence.
Table 16: Intrinsic OpenIE evaluation using the CaRB [6] framework. Performance difference between the 10 longest and 10 shortest annotated passages using our default GPT-3.5 Turbo (1106) model.
AUC Precision Recall F1
10 Shortest Passages $58.9$ $79.2$ $65.7$ $71.8$
10 Longest Passages $39.0$ $60.7$ $48.5$ $53.9$
@@ Appendix G Cost and Efficiency Comparison
Although offline indexing time and costs are higher for HippoRAG than IRCoT—around $10$ times slower and \$$15$ more expensive for every $10,000$ passages  [^7] To speed up indexing, we use $10$ threads querying gpt-3.5-turbo-1106 through the OpenAI API in parallel. At the time of writing, the cost of the API is \$$1$ for a million input tokens and \$$2$ for a million output tokens., these costs can be dramatically reduced by leveraging open source LLMs. As shown in our ablation study in Table 5 Llama-3.1-70B-Instruct [1] performs similarly to GPT-3.5 Turbo even though it can be deployed locally using vLLM [40] and 4 H100 GPUs to index 10,000 documents in around 4 hours, as seen in Table 18. Additionally, since these costs could be even further reduced by locally deploying this model, the barriers for using HippoRAG at scale could be well within the computational budget of many organizations. Finally, we note that even if LLM generation cost drops, the online retrieval efficiency gains discussed above remain intact given that the number of tokens required for IRCoT vs. HippoRAG stay constant and LLM use is likely to also remain the system’s main computational bottleneck.
Table 18: Average cost and latency measurements for offline indexing using GPT-3.5 Turbo and locally deployed Llama-3.1 (8B and 70B) using vLLM on $10,000$ passages.
@@
Llama-3.1-8B-Instruct API Cost (\$) 00 00 00
Time (minutes) $7$ $7$ $120$
Llama-3.1-70B-Instruct API Cost (\$) 00 00 00
Time (minutes) $7$ $7$ $250$
@@ Appendix H Implementation Details & Compute Requirements
In terms of compute requirements, most of our compute requirements are unfortunately not disclosed by the OpenAI. We run ColBERTv2 and Contriever for indexing and retrieval we use 4 NVIDIA RTX A6000 GPUs with 48GB of memory. For indexing with Llama-3.1 models, we use 4 NVIDIA H100 GPUs with 80GB of memory. Finally, we used 2 AMD EPYC 7513 32-Core Processors to run the Personalized PageRank algorithm.